# supReferences

### v4

@@ Abstract
[…]
@@ 1.1 Context
Denoting by $t\in\mathbb{N}$ the time index and by $\hat{\eta}_{t}$ the conditional distribution of the state variable at time $t$ given the available observations $\{y_{s}\}_{s=1}^{t}$, a PF is a Monte Carlo algorithm that propagates over time a set of $N\in\mathbb{N}$ particles that can be used to estimate, in an online fashion, the sequence of the so-called filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$. For instance, in self-driving cars applications $\hat{\eta}_{t}$ is the conditional distribution of the position of the car at time $t$ given its GPS localisation (known to have a precision of a few meters) and its distance with respect to some nearby landmarks, under the assumed SSM.
The popularity of PFs stems from the fact that this class of algorithms can be deployed on a very large class of SSMs. In addition, it is widely believed that the quality of the PF estimate $\hat{\eta}_{t}^{N}$ of $\hat{\eta}_{t}$ does not deteriorate over time, a belief supported by several theoretical analyses showing that, for some class of functions $\mathfrak{F}$ and $p\in\mathbb{N}$, we have (see Caffarel et al.,, 2025, and references therein)
@@ 1.2 Contributions of the paper
[…]
@@
For a given $\kappa\in(0,1)$ the PF estimate $(\hat{\eta}_{t}^{N})_{t\geq 1}$ of $(\hat{\eta}_{t})_{n\geq 1}$ is therefore necessarily such that we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$. Given the ever increasing use of PFs in life critical applications, notably due to the increasing prevalence of autonomous systems (Ogunsina et al.,, 2024) and the role that PFs play in these systems, it is then crucial to understand the behaviour of $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ as a function of the number of particle $N$ and of the time horizon $T$. This is particularly important for applications which require to accurately estimate a very long sequence of filtering distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$; for instance, according to Kalra and Paddock, (2016), in order to demonstrate the reliability of self-driving cars their onboard algorithms must be tested for hundreds of billions of miles.
@@
The second contribution of this paper is to provide a first sharp result on the dependence of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ to $N$ and $T$. Considering the same filtering problem as the one described above, we prove that there exists a finite constant $C_{\kappa}$ such that for any $q\in(0,1)$ and time horizon $T\geq 1$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N_{\kappa,T,q}:=C_{\kappa}(\log(T)+\log(1/q))$. For a given $\kappa$, the dependence of $N_{\kappa,T,q}$ to $T$ and $q$ is shown to be sharp, and thus to guarantee that we still have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ as $T$ increases the number of particles only needs to grow at speed $\log(T)$ while, to decrease $q$, the number of particles should only grow at speed $\log(1/q)$. It is worth noting at this stage that Corollary 14.5.7 in Del Moral, (2013) establishes that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N^{\prime}_{\kappa,T,q}:=C^{\prime}_{\kappa}(\log(T+e)(1+\log(1/q))$ and for some finite constant $C^{\prime}_{\kappa}$. We also note that it has not been realized (and therefore proved) in this reference that $N$ must increase with $T$ to ensure that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)$ remains bounded by $q$ as $T$ increases (and thus that for any fixed $N\geq 1$ we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$). We finally remark that this result is obtained under (strong) assumptions which are not verified for the filtering problem that we consider in this work.
@@
Despite the reassuring results we proved regarding the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often. Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte (RQMC) point sets. This idea has been proposed and studied in Gerber and Chopin, (2015) with the aim of obtaining a filtering algorithm, called sequential quasi-Monte Carlo (SQMC), which has a faster convergence rate (as $N\rightarrow\infty$) than PFs. Letting $(\tilde{\eta}^{N}_{t})_{t\geq 1}$ denote the estimate of $(\hat{\eta}_{t})_{t\geq 1}$ obtained with SQMC, the third contribution of this paper is to show, for the one-dimensional filtering problem outlined above, $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\tilde{\eta}^{N}_{t}-\hat{\eta}_{t}\|=0$ with probability one.
To analyse the time evolution of PF and SQMC estimates we introduce a new approach for studying PF algorithms which builds on tools developed in the PF and in the quasi-Monte Carlo literature. Informally speaking, in a first step we analyse the behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ for a given realization of the $\mathcal{U}(0,1)$ random numbers underpinning the algorithm. Then, in a second step, we leverage the known properties of sets of independent $\mathcal{U}(0,1)$ random numbers and of RQMC point sets to apply the result obtained in the first step to PFs and SQMC.
@@ 1.3 Notation and organization of the paper
In what follows all the random variables are defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. To introduce some notation let $s\in\mathbb{N}$. Then, we let $\mathcal{B}(\mathbb{R}^{s})$ denote the Borel $\sigma$-algebra on $\mathbb{R}^{s}$, $\mathcal{P}(\mathbb{R}^{s})$ denote the set of probability measures on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, $\lambda_{s}$ denote the Lebesgue measure on $\mathbb{R}^{s}$ and
$\mathcal{B}_{s}=\Big{\{}B\subset\mathbb{R}^{s}:B=\prod_{i=1}^{s}(-\infty,a_{i}],\,(a_{1},\dots a_{s})\in\mathbb{R}^{s}\Big{\}}.$
In addition, for any signed-measure $\mu$ on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$ we let $\|\mu\|=\sup_{B\in\mathcal{B}_{s}}|\mu(B)|$ and, for a point set $\{u^{n}\}_{n=1}^{N}$ in $[0,1)^{s}$, we denote by $D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}$ its star discrepancy, that is
$D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}:=\sup_{b\in[0,1)^{s}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{[0,b)}(u^{n})-\lambda_{s}\big{(}[0,b)\big{)}\Big{|}.$
@@
Next, for a probability distribution $\pi\in\mathcal{P}(\mathbb{R}^{s})$, a measurable function $f:\mathbb{R}^{s}\rightarrow\mathbb{R}$ and a (potentially un-normalized) kernel $K$ acting on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, we let $\pi(f)=\int_{\mathbb{R}^{s}}f(x)\pi(\mathrm{d}x)$, $\pi K$ denote the measure such that $\pi K(A)=\int_{\mathbb{R}^{s}}K(x,A)\pi(\mathrm{d}x)$ for all $A\in\mathcal{B}(\mathbb{R}^{s})$, and $(K^{k})_{k\geq 1}$ denote the sequence of (potentially un-normalized) kernels acting on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$ defined by
$K^{k}(x,A)=\int_{\mathbb{R}^{s}}K(x,\mathrm{d}x^{\prime})K^{k-1}(x^{\prime},A),\quad A\in\mathcal{B}(\mathbb{R}^{s}),\quad k\geq 1$
with the convention that $K^{k-1}(x^{\prime},\mathrm{d}x)=\delta_{\{x^{\prime}\}}(\mathrm{d}x)$ for all $x^{\prime}\in\mathbb{R}^{s}$ when $k=1$. Finally, for any probability distribution $\pi\in\mathcal{P}(\mathbb{R})$ we let $F_{\pi}$ denote the cumulative density function of $\pi$ and $F^{-1}_{\pi}$ denote its (generalized-)inverse.
The rest of the paper is organized as follows. Section 2 introduces formally the filtering problem and algorithms discussed in this work, and the time evolution of the resulting PF estimates is studied in Section 3. The result outlined in the introductory section for SQMC is given in Section 4 and Section 5 concludes. All the proof are gathered in the appendix.

### v5

# Safety of particle filters: Some results on the time evolution of particle filter estimates

@@ Abstract
Particle filters (PFs) form a class of Monte Carlo algorithms that propagate over time a set of $N\geq 1$ particles which can be used to estimate, in an online fashion, the sequence of filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$ defined by a state-space model. Despite the popularity of PFs, the study of the time evolution of their estimates has received barely any attention in the literature. Denoting by $(\hat{\eta}_{t}^{N})_{t\geq 1}$ the PF estimate of $(\hat{\eta}_{t})_{t\geq 1}$ and letting $\kappa\in(0,1/2)$, in this work we first show that for any number of particles $N$ it holds that, with probability one, we have $\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa$ for infinitely many time instants $t\geq 1$, with $\|\cdot\|$ the Kolmogorov distance between probability distributions. Considering a simple filtering problem we then provide reassuring results concerning the ability of PFs to estimate jointly a finite set $\{\hat{\eta}_{t}\}_{t=1}^{T}$ of filtering distributions by studying the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$. Finally, on the same toy filtering problem, we prove that sequential quasi-Monte Carlo, a randomized quasi-Monte Carlo version of PF algorithms, offers greater safety guarantees than PFs in the sense that, for this algorithm, it holds that $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=0$ with probability one.
@@ 1.1 Context
Denoting by $t\in\mathbb{N}:=\{1,2,\dots\}$ the time index and by $\hat{\eta}_{t}$ the conditional distribution of the state variable at time $t$ given the available observations $\{y_{s}\}_{s=1}^{t}$, a PF is a Monte Carlo algorithm that propagates over time a set of $N\in\mathbb{N}$ particles that can be used to estimate, in an online fashion, the sequence of the so-called filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$. For instance, in self-driving cars applications $\hat{\eta}_{t}$ is the conditional distribution of the position of the car at time $t$ given its GPS localisation (known to have a precision of a few meters) and its distance with respect to some nearby landmarks, under the assumed SSM.
The popularity of PFs stems from the fact that this class of algorithms can be deployed on a very large class of SSMs. In addition, it is widely believed that the quality of the PF estimate $\hat{\eta}_{t}^{N}$ of $\hat{\eta}_{t}$ does not deteriorate over time, a belief supported by several theoretical analyses showing that, for some class of functions $\mathfrak{F}$ and constant $p\in\mathbb{N}$, we have (see Caffarel et al.,, 2025, and references therein)
@@ 1.2 Contributions of the paper
In this paper we first show that such a theoretical guarantee for PFs cannot exist. Formally, denoting by $\|\cdot\|$ the Kolmogorov distance between probability distributions, we prove that for any $\kappa\in(0,1/2)$, with probability one and for any number of particles $N\geq 1$, we have $\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa$ for infinitely many time instants $t\geq 1$. The fact that in some challenging situations a PF may fail to accurately estimate some filtering distributions is not surprising: it is well-known that this problem will usually happen when the distance between two successive filtering distributions $\hat{\eta}_{t}$ and $\hat{\eta}_{t+1}$ is large, e.g. because the SSM is not time homogenous or because the observations $y_{t}$ and $y_{t+1}$ are very different from each other. However, and crucially, we prove the above property of PFs on what is arguably the simplest filtering scenario, namely for a one-dimensional and time-homogenous linear Gaussian SSM with observations $y_{t}=0$ for all $t\geq 1$. As we will see in what follows, this negative result for PF estimates is trivial to prove, and arises from the very simple and intuitive reason that too much randomness is used to propagate the particle system over time. We stress that this problem is not limited to PFs but is common and intrinsic to any sequential (importance) sampling algorithms, including the ensemble Kalman filter (Roth et al.,, 2017).
@@
For any given $\kappa\in(0,1)$, the PF estimate $(\hat{\eta}_{t}^{N})_{t\geq 1}$ of the sequence $(\hat{\eta}_{t})_{n\geq 1}$ is therefore necessarily such that $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$. Given the ever-increasing use of PFs in life critical applications, notably due to the increasing prevalence of autonomous systems (Ogunsina et al.,, 2024) and the role that PFs play in these systems, it is then crucial to understand the behaviour of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ as a function of the number of particles $N$ and of the time horizon $T$.
The second contribution of this paper is to provide a first sharp result on the dependence of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ to $N$ and $T$. Considering the same filtering problem as the one described above, we prove that there exists a constant $C\in(0,\infty)$ such that, for any $(\kappa,q)\in(0,1)^{2}$ and any $T\geq 1$, the inequality $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ holds for all $N\geq N_{T,\kappa,q}:=Cv_{\kappa}\log(T/q)$, with $v_{\kappa}=\{1+\log(1+\kappa^{-1})\}^{2}\kappa^{-2}$. We also establish that, for a given $\kappa$, the dependence of $N_{T,\kappa,q}$ on the two parameters $T$ and $q$ is sharp.
@@
To the best of our knowledge, prior to this work the only study of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ was given in Del Moral, (2013, Corollary 14.5.7, page 449). In this reference, it is shown that the inequality $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ holds for all $N\geq N^{\prime}_{T,\kappa,q}:=C^{\prime}\kappa^{-2}\log(1+T)(1+\log(1/q))$, with $C^{\prime}$ a finite constant. This result is however obtained under strong assumptions, which are for instance not satisfied for the filtering problem that we consider in this work. We also note that $N^{\prime}_{T,\kappa,q}$ has worse dependence on the parameters $(T,q)$ than $N_{T,\kappa,q}$, although its dependence on $\kappa$ is slightly better. Lastly, and importantly, it has not been realized (and therefore proved) in Del Moral, (2013) that $N$ must increase with $T$ to ensure that the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)$ remains bounded above by $q$ as $T$ increases, and thus that for any fixed $N\geq 1$ we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$.
@@
Despite the reassuring results we proved regarding the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often. Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte point sets. This idea has been proposed and studied in Gerber and Chopin, (2015) with the aim of obtaining a filtering algorithm, called sequential quasi-Monte Carlo (SQMC), which has a faster convergence rate (as $N\rightarrow\infty$) than PFs. Letting $(\tilde{\eta}^{N}_{t})_{t\geq 1}$ denote the estimate of $(\hat{\eta}_{t})_{t\geq 1}$ obtained with SQMC, the third contribution of this paper is to show, for the one-dimensional filtering problem outlined above, that we have $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\tilde{\eta}^{N}_{t}-\hat{\eta}_{t}\|=0$ with probability one.
@@ 1.3 Notation and organization of the paper
In what follows all the random variables are defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$, and for any $s\in\mathbb{N}$ we let $\mathcal{B}(\mathbb{R}^{s})$ denote the Borel $\sigma$-algebra on $\mathbb{R}^{s}$, $\mathcal{P}(\mathbb{R}^{s})$ denote the set of probability measures on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, $\lambda_{s}$ denote the Lebesgue measure on $\mathbb{R}^{s}$, and for any point set $\{u^{n}\}_{n=1}^{N}$ in $0,1)^{s}$ we denote by $D^{*}_{s}\big(\{u^{n}\}_{n=1}^{N}\big)$ its star discrepancy, that is
$D^{*}_{s}\big(\{u^{n}\}_{n=1}^{N}\big)=\sup_{b\in[0,1)^{s}}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{[0,b)}(u^{n})-\lambda_{s}\big([0,b)\big)\Big|.$
Next, for any probability measure $\pi\in\mathcal{P}(\mathbb{R})$ we let $F_{\pi}$ denote its cumulative density function and $F^{-1}_{\pi}$ denote the (generalized-)inverse of $F_{\pi}$. In this notion, the Kolmogorov distance between the probability measures $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R})$ is given by $\|\pi_{1}-\pi_{2}\|=\sup_{a\in\mathbb{R}}|F_{\pi_{1}}(a)-F_{\pi_{2}}(a)|$, and we denote by $\|\pi_{1}-\pi_{2}\|_{\mathrm{D}}$ the discrepancy metric between these two probability measures, that is we let
$\|\pi_{1}-\pi_{2}\|_{\mathrm{D}}=\sup_{-\infty<a<b<\infty}\big|\pi_{1}([a,b])-\pi_{2}([a,b])\big|.$
We recall the reader that these two norms on $\mathcal{P}(\mathbb{R})$ are related together through the inequalities
$\|\pi_{1}-\pi_{2}\|\leq\|\pi_{1}-\pi_{2}\|_{\mathrm{D}}\leq 2\|\pi_{1}-\pi_{2}\|,\quad\forall\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}).$ (2)
With the exception of Propositions 2 and 3, the results presented below are expressed in term of the Kolmogorov metric, which is arguably more popular than the discrepancy metric. These results can however be easily expressed in term of the latter metric by using (2). Propositions 2 and 3 are expressed using the discrepancy metric for reasons that will be provided in due course (see Remarks 2 and 3).
@@
To introduce some further notation we let $\pi\in\mathcal{P}(\mathbb{R})$, $f:\mathbb{R}\rightarrow\mathbb{R}$ be a measurable function and $K$ be a (potentially un-normalized) kernel acting on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. Then, we let $\pi(f)=\int_{\mathbb{R}}f(x)\pi(\mathrm{d}x)$, $\pi K$ denote the measure on $\mathcal{B}(\mathbb{R})$ such that $\pi K(A)=\int_{\mathbb{R}}K(x,A)\pi(\mathrm{d}x)$ for all $A\in\mathcal{B}(\mathbb{R})$, and $(K^{k})_{k\geq 1}$ denote the sequence of (potentially un-normalized) kernels acting on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$ defined by
$K^{k}(x,A)=\int_{\mathbb{R}}K(x,\mathrm{d}x^{\prime})K^{k-1}(x^{\prime},A),\quad(x,A)\in\mathbb{R}\times\mathcal{B}(\mathbb{R}),\quad k\geq 1.$
Lastly, below we let $\Phi(\cdot)$ denote the cumulative density function of the $\mathcal{N}(0,1)$ distribution.
The rest of the paper is organized as follows. In the next subsection we provide a brief reminder on scrambled $(t,s)$-sequences. Section 2 introduces formally the filtering problem and algorithms studied in this work, and our main results are presented in Section 3. Main results not proved in Section 3 are proved in Section 4, while Section 5 contains the proof of Proposition 1 presented Section 2 as well as the proof of Lemma 1 stated Section 4, which is the main technical contribution of the paper.
1.4 Reminder on scrambled $(t,s)$-sequences
Let $t\geq 0$, $s\geq 1$ and $b\geq 2$ be three integers, and let
be the set of all $b$-ary boxes. Then, for an integer $m\geq t$ the point set $\{u^{n}\}_{n=1}^{b^{m}}$ is called a $(t,m,s)$-net in base $b$ if every $b$-ary box of volume $b^{t-m}$ contains exactly $b^{t}$ points, while a sequence $(u^{n})_{n\geq 1}$ of points in $0,1)^{s}$ is called a $(t,s)$-sequence in base $b$ if, for any integers $a\geq 0$ and $m\geq t$, the point set $\{u^{n}\}_{n=ab^{m}+1}^{(a+1)b^{m}}$ is a $(t,m,s)$-net in base $b$.
From the above definitions it should be clear that, from a practical perspective, $(t,s)$-sequences in base $b$ with $t=0$ and $b=2$ are preferable, but such sequences exist only for $s\in\{1,2\}$ (Niederreiter,, 1992, Corollary 4.24, page 62). In the context of this work, the crucial property of a $(t,s)$-sequence $(u^{n})_{n\geq 1}$ in base $b$ is that, for some constant $C_{s,t,b}\in(0,\infty)$ depending only on $s$, $t$ and $b$, we have (see Niederreiter,, 1992, Section 4.1)
Finally, a sequence $(U^{n})_{n\geq 1}$ of $(0,1)^{s}$-valued random variables is a scrambled $(t,s)$-sequence in base $b$ if, $\mathbb{P}$-a.s., $(U^{n})_{n\geq 1}$ is a $(t,s)$-sequence in base $b$ and if $U^{n}\sim\mathcal{U}(0,1)^{s}$ for all $n\geq 1$. Consequently, a scrambled $(t,s)$-sequence in base $b$ is a sequence $(U^{n})_{n\geq 1}$ of dependent $\mathcal{U}(0,1)^{s}$ random variables such that $D^{*}_{s}\big(\{U^{n}\}_{n=1}^{N}\big)\leq C_{s,t,b}N^{-1}\log(N)^{s}$ with $\mathbb{P}$-probability one and for all $N\geq 1$, and with the constant $C_{s,t,b}$ as in (3). A scrambled $(t,s)$-sequence is obtained by randomizing, or scrambling, a $(t,s)$-sequence, and various scrambling methods have been proposed in the literature. It is worth mentioning that the first $N$ points of a scrambled sequence can be generated in $\mathcal{O}(N\log(N))$ operations, but the memory requirements of the algorithm needed to perform this task depend on the scrambling procedure that is used (Hong and Hickernell,, 2003). We refer the reader to Chapter 3 in Dick and Pillichshammer, (2010) and to Owen, (2023) for a discussion on scrambling algorithms.